OpenStudy (anonymous):

A man on the tenth floor of a building shouts down to a person on the street. If the man is 25 feet up and the distance between the person on the street and the man in the building is 50 feet, what is the angle of elevation (in degrees) between the person on the street and the person in the building? A) 15° B) 30° C) 45° D) 60°

OpenStudy (cwrw238):



OpenStudy (cwrw238):

sin x = 25/50 = 1/2 angle whose sine is 1/2 = x = 30 degrees so B